ebjohnson1:

5 - 2x2 = -15 solve it using inverse operations

WilliamMiller08:

5 - 2x^2 - 5 = -15 - 5 -2x^2 = -20 -2x^2/-2 = -20/-2 x^2=10

WilliamMiller08:

I think this is right.